Suppose that the probability that a person books an airline ticket using an online travel website is 0.79. For the question that follows, consider a sample of ten randomly selected people who recently booked an airline ticket.

1. What is the probability that exactly eight out of ten people used an online travel website when they booked their airline ticket?

2. What is the probability that at least nine out of ten people used an online travel website when they booked their airline ticket?
Answer #1

Solution - m -X xn Binomial (10; 0.79) 10 p = 0.79 Pix) = (?) bta max 8 10-8 )) D) Plx = 8) = 10) (0-19) (0-2)) 10.3011 9 10-